# A Formal, Symbolic Analysis of the Matrix Cryptographic Protocol Suite

### v1

@@ Abstract
Using Verifpal, we construct formal models of Olm and Megolm, as well as their composition. We prove various properties of interest about Olm and Megolm, including authentication, confidentiality, forward secrecy, and post-compromise security. We also mechanize known limitations, previously discovered attacks, and trivial attacker wins from the specifications and previous literature. Finally, we model Sender Keys and the composition of Signal with Sender Keys in order to draw a comparison with Olm, Megolm, and their composition. Our mechanized analysis indicates subtle yet critical differences in forward secrecy between Olm and Signal, which in turn affects the post-compromise security of Megolm and Sender Keys, respectively. From our analysis we conclude the composition of Olm and Megolm has comparable security to the composition of Signal and Sender Keys if Olm pre-keys are signed, and provably worse post-compromise security if Olm pre-keys are not signed.
@@ I Introduction
Forward secrecy and post-compromise security are very well understood for secure two-peer messaging protocols such as Signal [39]. However, secure group messaging has comparatively seen much less attention [96]. Group messaging introduces additional complexity, as members join and leave over time, motivating novel protocol designs. One notable protocol for secure instant group messaging is the Sender Keys variant of Signal [72] which, among other applications, is employed by WhatsApp [101], Facebook Messenger [51], the Signal Application [89], and the Session Application [87].
Matrix is a recently introduced open standard for secure, decentralized, real-time communication that promises interoperability and end-to-end encryption [37]. Unlike Signal, Matrix is designed to be deployed in a federated way, giving users control over which servers host and manage their conversations. Matrix has seen wide adoption across governments, the private sector, and the general public. It is used by government organizations in France, Germany, Sweden, and Luxembourg; for example, France’s central administration takes place on a Matrix-based internal network, Tchap [50], and Germany’s defense and healthcare organizations employ Matrix in the field [55]. Mozilla and KDE use Matrix internally [41]. A number of existing applications integrate Matrix, including the forum software Discourse [85] and the enterprise messaging platform Rocket.Chat [85]. As of April 2024 Matrix reports over 115 million users on public, data-reporting home servers [95].
@@
Our contribution. In this work we take an approach rooted in formal methods and computer-aided reasoning to study the cryptographic security of the Matrix protocol suite, allowing us to construct computer-verified proofs or verified and explicit counterexamples. We employ Verifpal as our symbolic analysis tool of choice. We selected Verifpal, as opposed to ProVerif or Tamarin, because it is simpler and thus better suited for comparative analysis. It also provides easy extractions to Proverif and Rocq111Formerly known as Coq for more precise verification if necessary [65]. We make the following contributions:
Models. We present the first symbolic analysis and mechanization of the Matrix protocol suite; taking such an approach allows us to precisely and automatically compare the cryptographic guarantees of the Matrix cryptographic suite to its contemporaries. We construct the first formal models for Olm, Megolm, and their composition. To understand how Olm and Megolm compare with Signal and Sender Keys, we employ a pre-existing Verifpal Signal model [65], we implement Sender Keys, and we model the composition of Signal and Sender Keys.
Verification. Using Verifpal for automated symbolic analysis, we derive various proofs, limitations, and previously found attacks from the Matrix and Sender Keys specifications and previous formalizations. First, we mechanically prove the confidentiality, authentication, forward secrecy, and post-compromise security properties of Olm and Megolm, mechanizing the previous hand-written computational analysis of the Matrix cryptographic core [11] and the stated security goals in the Olm [53] and Megolm [52] specifications. We then mechanize previously known attacks and limitations, including the Megolm unknown key-share attack described in the Matrix NCC audit [58], the tradeoff between forward secrecy and deniability in Olm [53], and the unauthorized injection attack described in the computational analysis of Sender Keys [17]. We mechanically prove the extension of Sender Keys proposed by Balbas et al. satisfies the intended security properties, confidentiality and authentication [17].
Comparison. Using our Verifpal models, we precisely compare the cryptographic guarantees of Olm and Megolm with Signal and Sender keys. Among other insights, we observe Olm channels without pre-key signing result in the post-compromise security of the composition of Megolm being significantly weaker than Sender Keys. We note that the Megolm specification [52] does not require Olm pre-keys to be signed while the Matrix protocol [2] requires the Olm pre-keys to be signed but provides no reason why. Our work elucidates this question providing formal proofs that in the absence of Olm pre-keys signed, the composition of Olm and Megolm provides weaker security.
@@ II-A Overview of Signal and Sender Keys
The Signal handshake is defined in Figure 6 in the Appendix, and we refer the reader to [39] for a complete formal description of the protocol.
@@ II-B Overview of Olm and Megolm
Megolm. Megolm is used to ensure group communications are encrypted. Originally introduced in 2016, before Sender Key (group messaging variant of Signal) [72] was published, Megolm has very similar functionality to Sender Keys. However, Megolm utilizes Olm as opposed to Signal as the secure peer-to-peer channel in order to share Megolm session information between peers. Additionally, Megolm utilizes a custom ratchet scheme that can be advanced an arbitrary amount forward while needing at most 1020 hash computations [52], which is further described in the Appendix VIII-A. Aside from the details in cryptographic primitives between implementations of Sender Keys and Megolm as shown in Table II, the secure channel protocol, and the ratcheting scheme, all other details between the two protocols remain the same.
@@ III Olm and Megolm Formal Models
In this section we overview the Verifpal automated symbolic reasoning tool, the subsequent abstractions and limitations of our models, the details of our models, and the properties of interest we seek to reason about.
III-A Modeling in Verifpal
To assess the cryptographic properties of Olm and Megolm, we employ Verifpal, a cryptographic protocol verification tool for the symbolic (Dolev-Yao) model. In stark comparison to the hardly accessible heavyweight tools in the space, Tamarin [20] and ProVerif, Verifpal seeks to strike a balance between intuitive and simple protocol modeling whilst still being precise and providing useful results [65]. Verifpal has been previously used to analyze Signal [65], DP3T [65], Zoom’s cryptographic suite [74], Ockam’s cryptographic protocol suite [76], and post-quantum protocols for cyber-physical systems [79]. We also observe analyzing Olm, Megolm, Signal, and Sender Keys (in the symbolic model) does not require any cryptographic primitives or setups outside the expressibility or analysis capability of Verifpal. Thus we deem Verifpal an effective fit for our analysis goals. We note that while Verifpal is sufficient for our analysis, if we sought to analyze a stateful protocol, required more complex primitives, or needed more efficient and refined proof methods, Tamarin or ProVerif would have been a better fit.
Protocols in Verifpal are modeled using a formal language that specifies the steps and messages exchanged between participants. Participants can be specified to have pre-specified knowledge of certain values, generate new values, apply various (predefined) cryptographic primitives to values, leak values, or send values to other participants. Queries can then be specified to assess, for example, the confidentiality or authenticity of a given value. The attacker can either be specified as active, indicating full control over the channel, or passive, indicating only observational ability over the channel. To illustrate the modeling scheme of Verifpal, see the Triple Diffie-Hellman example in Figure 1.
To ensure Verifpal terminates and is always able to complete a proof (or discover a counterexample)222Recall reasoning about confidentiality, authentication, post-compromise security, and forward secrecy under Dolev-Yao is computationally decidable, so we know Verifpal will always eventually terminate given enough resources., we ran our experiments on a research computing cluster with 88 cascade lake CPUs and 5TB of RAM.
@@ III-B Model Details
Olm. We model the complete cryptographic handshake of Olm. The semantics of Verifpal aligns very nicely to the semantics of Olm, allowing us to create a very precise model. A diagram of the modeled handshake is shown in Appendix Section 2.
Megolm. We model the initialization of a Megolm session, the transmission of the ratchet key to peers in the Megolm session via Olm, the exchange of messages between peers in a Megolm session, and the incrementation of the Megolm ratchet. A diagram of the modeled cryptographic exchange is shown in Appendix Section 3.
Signal. We employ a previously created Signal model, as presented in the Verifpal manual [65].
Sender Keys. We model the initialization of a Sender Keys session, the transmission of the ratchet key to peers in the Sender Keys session via Signal, the exchange of messages between peers in a Sender Keys session, and the incrementation of the Sender Keys ratchet. A diagram of the modeled cryptographic exchange is shown in Appendix Section 3.
@@ III-C Abstractions and Limitations
Perfect Underlying Primitives. As in the symbolic cryptanalysis model, we assume the underlying cryptographic primitives are black-boxes. That is, we assume the primitives used in our models are secure and correct, particularly HKDF, HMAC, AEAD, and key signing. Verifpal and other cryptographic reasoning in tools in the symbolic model traditionally assume perfect cryptography.
No Ratchet Reseeding. The Megolm ratchet is designed to re-seed after every iterations as described in the Appendix VIII-A. We do not capture this reseeding in our model — only forward iterations — as it is not relevant to the analysis under the Dolev-Yao model.
Finite Peers. We reason about a finite number of peers in our model, and prove our properties of interest under that assumption.
@@ IV-A Proven Properties of Interest
Using our Verifpal models, we automatically prove the Olm and Megolm properties as stated and motivated in section III-D. This includes Olm message confidentiality and authentication, Olm session key confidentiality, Olm forward secrecy (assuming Olm chooses to sign pre-keys), Olm post-compromise security, Megolm message confidentiality and authentication, and Megolm forward secrecy.
@@ V-A Computer-Aided Protocol Verification
In comparison, manual symbolic cryptanalysis has been theoretically feasible [29] but seldom employed until recent advances in proof-oriented programming languages. In particular, F*, a general-purpose proof-oriented language with support for dependent types proving, refinement times, executable semantics extractions, and verified implementation extractions to F#, C, or Wasm [94], has seen usage for this purpose. Additionally, other dependent types-based theorem provers such as Rocq [24] and Agda [75] are suited for this purpose, as they both offer executable semantics and implementation extractions 333Note, Rocq and Agda have different type theories: Rocq is based on the Calculus of Constructions (CoC) while Agda is based on Martin-Löf Type Theory. This difference results in different approaches to constructing proofs. It is entirely unexplored how this difference affects the formulation of cryptographic protocols via the methodology described in [29].. The theoretical framework for verifying protocols symbolically using dependent-types introduced in [29] has been formalized in F* and was used to analyze Signal [27] and later the Message Layer Security protocol [26].444Note that F* and dependent-types theorem provers are incredibly general and are expressive enough to capture all of mathematics. Interestingly, F* has been additionally used to construct verified cryptographic libraries [105] and to reason about group key agreement [25]. However, constructing such formalizations in a dependent types prover such as F* is generally much more time consuming than constructing models in a symbolic cryptanalysis tool.
@@
Distributed Protocol Analysis. For verifying non-cryptographic properties of distributed protocols such as safety, liveness, or absence of deadlocks, a separate body of work utilizing a different flavor of computer-aided tools exists. In general, resolving safety, liveness, and deadlock-free properties reduces to a PSPACE-complete automata-theoretic decision procedure: Büchi Automata language inclusion [82]. Similarly, there is a long line of work specifying distributed systems in Hoare-style program logics and automatically reasoning about them [86] [88]. Most similarly to our work, Model checkers have been used automatically (and decidably) reason about properties of specified distributed protocols. Although model checking is generalizable for reasoning about general programs [38], TLA+ and Spin are specialized model checkers for reasoning about distributed and concurrent systems [67] [60]. These model checkers have been used to analyze Paxos [45] [68], Raft [77], Chord [103], the Lightning Network [100], TCP [78] [99], SCTP [56], and DCCP [78], among many other protocols. For systems with more nuanced properties such as peer behavior scoring in Ethereum’s GossipSub [66] and RTT calculations for TCP packets [98], constructive, manual proof assistants such as ACL2 [61] have been employed.
@@ VI Conclusion
This paper has presented a comprehensive formal analysis of the Olm and Megolm protocols within the Matrix cryptographic suite. Utilizing Verifpal, we constructed precise models and conducted automated symbolic analysis to verify key security properties, including confidentiality, authentication, forward secrecy, and post-compromise security for Olm and Megolm. Our approach enabled us to precisely and automatically reproduce known vulnerabilities and attacks. In addition, through our comparative analysis with the Signal and Sender Keys protocols, we highlighted critical differences in post-compromise security, emphasizing the need for ongoing scrutiny and improvement of secure messaging protocols. We endorse the Matrix protocol specification’s current decision for Olm pre-keys to be signed, as our models confirm that when Olm pre-keys are signed the security guarantees of Olm and Megolm are comparable to Signal and Sender Keys. Our findings underscore the importance of formal, mechanized proofs in ensuring the robustness of cryptographic systems, particularly in the context of federated and decentralized communication platforms like Matrix.
VII Acknowledgements
Jacob Ginesin would like to thank Zack Eisbach, Thomas Dwyer, and Eddie Berman for many insightful discussions about formal methods and protocol modeling.
@@ VII Acknowledgements
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.4cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.0cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.2cm ]
[…]
[4.0]
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.4cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.2cm ]
[…]
[6.75]
[…]
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.4cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.2cm ]
[…]
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.4cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.2cm ]
[…]

### v2

# The Matrix Reloaded: A Mechanized Formal Analysis of the Matrix Cryptographic Suite

@@ Abstract
Using ProVerif, we construct formal models of Olm and Megolm, as well as their composition. We prove various properties of interest about Olm and Megolm, including authentication, confidentiality, forward secrecy, and post-compromise security. We also mechanize known limitations, previously discovered attacks, and trivial attacker wins from the specifications and previous literature. Finally, we model Sender Keys and the composition of Signal with Sender Keys in order to draw a comparison with Olm, Megolm, and their composition. Our mechanized analysis indicates subtle yet critical differences in forward secrecy between Olm and Signal, which in turn affects the post-compromise security of Megolm and Sender Keys respectively. From our analysis we conclude the composition of Olm and Megolm has comparable security to the composition of Signal and Sender Keys if Olm pre-keys are signed, and provably worse post-compromise security if Olm pre-keys are not signed.
@@ I Introduction
Forward secrecy and post-compromise security are well understood for secure two-peer messaging protocols such as Signal [39]. However, secure group messaging has comparatively seen much less attention [83]. Group messaging introduces additional complexity, as members join and leave over time, motivating novel protocol designs. One notable protocol for secure instant group messaging is the Sender Keys variant of Signal [67] which, among other applications, is employed by WhatsApp [85], Facebook Messenger [50], the Signal Application [76], and the Session Application [75].
Matrix is a recently introduced open standard for secure, decentralized, real-time communication that promises interoperability and end-to-end encryption [38]. Unlike Signal, Matrix is designed to be deployed in a federated way, giving users control over which servers host and manage their conversations. Matrix has seen wide adoption across governments, the private sector, and the general public. It is used by government organizations in France, Germany, Sweden, and Luxembourg; for example, France’s central administration takes place on a Matrix-based internal network, Tchap [49], and Germany’s defense and healthcare organizations employ Matrix in the field [56]. Mozilla and KDE use Matrix internally [41]. A number of existing applications integrate Matrix, including the forum software Discourse [74] and the enterprise messaging platform Rocket.Chat [74]. As of September 2024 Matrix reports over 115 million users on public, data-reporting home servers [82].
@@
Our contribution. In this work we take an approach rooted in formal methods and computer-aided reasoning to study the cryptographic security of the Matrix protocol suite, allowing us to construct computer-verified proofs or verified and explicit counterexamples. We employ ProVerif as our symbolic analysis tools of choice. We make the following contributions:
Models. We present the first symbolic analysis and mechanization of the Matrix protocol suite; taking such an approach allows us to precisely and automatically compare the cryptographic guarantees of the Matrix cryptographic suite to its contemporaries. We construct the first formal models for Olm, Megolm, and their composition. To understand how Olm and Megolm compare with Signal and Sender Keys, we employ a pre-existing ProVerif Signal model [62], we implement Sender Keys, and we model the composition of Signal and Sender Keys.
Verification. Using ProVerif for automated symbolic analysis, we derive various proofs, limitations, and previously found attacks from the Matrix and Sender Keys specifications and previous formalizations. First, we mechanically prove the confidentiality, authentication, forward secrecy, and post-compromise security properties of Olm and Megolm, mechanizing the previous hand-written computational analysis of the Matrix cryptographic core [11] and the stated security goals in the Olm [52] and Megolm [51] specifications. We then mechanize previously known attacks and limitations, including the Megolm unknown key-share attack described in the Matrix NCC audit [58], the tradeoff between forward secrecy and deniability in Olm [52], and the unauthorized injection attack described in the computational analysis of Sender Keys [17]. We mechanically prove the extension of Sender Keys proposed by Balbas et al. satisfies the intended security properties, confidentiality and authentication [17].
Comparison. Using our ProVerif models, we precisely compare the cryptographic guarantees of Olm and Megolm with Signal and Sender keys. Among other insights, we observe Olm channels without pre-key signing result in the post-compromise security of the composition of Megolm being significantly weaker than Sender Keys. We note that the Megolm specification [51] does not require Olm pre-keys to be signed while the Matrix protocol [2] requires the Olm pre-keys to be signed but provides no reason why. Our work elucidates this question providing formal proofs that in the absence of Olm pre-keys signed, the composition of Olm and Megolm provides weaker security.
Ethics consideration. We have alerted the Matrix foundation’s security contact point of our findings and risks. They pointed us to the design document specifiying the Matrix architecture [2] that require Olm pre-keys to be pre-signed, thus confirming our findings.
@@ II-A Overview of Signal and Sender Keys
The Signal handshake is visualized in Figure 5 in the Appendix, and we refer the reader to [39] for a complete formal description of the protocol.
@@ II-B Overview of Olm and Megolm
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.3cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.1cm ]
[…]
@@
Megolm. Megolm is used to ensure group communications are encrypted. Originally introduced in 2016, before Sender Key (group messaging variant of Signal) [67] was published, Megolm has very similar functionality to Sender Keys. However, Megolm utilizes Olm as opposed to Signal as the secure peer-to-peer channel in order to share Megolm session information between peers. Additionally, Megolm utilizes a custom ratchet scheme that can be advanced an arbitrary amount forward while needing at most 1020 hash computations [51], which is further described in Figure VII-A. Aside from the details in cryptographic primitives between implementations of Sender Keys and Megolm as shown in Table II, the secure channel protocol, and the ratcheting scheme, all other details between the two protocols remain the same.
@@ II-D Scope of Our Analysis
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.3cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.0cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.1cm ]
[…]
@@ III Olm and Megolm Formal Models
We fully model the Matrix protocol suite as well as Sender Keys using ProVerif, an automated symbolic analysis tool in the Dolev-Yao model. In this section we overview the ProVerif automated symbolic reasoning tool, the subsequent abstractions and limitations of our models, the details of our models, and the properties of interest we seek to reason about.
III-A Modeling in ProVerif
We choose ProVerif as our symbolic protocol analysis tool of choice, boasting an efficient pi-calculus-based proof engine which reduces the term rewriting search problem that characterizes symbolic cryptanalysis to horn clauses, as well as an expressive syntax for specifying protocols [34].
To model protocols in ProVerif, we begin by defining each participant as a process, representing the roles they play within the protocol. Each process consists of message exchanges, including cryptographic operations such as encryption, decryption, signing, and verification. Cryptographic primitives are modeled symbolically using predefined functions in ProVerif, which capture their idealized behavior without considering their computational implementation.
The adversary is explicitly modeled with the ability to intercept, modify, and inject messages into the communication channel. However, they are restricted by the symbolic security assumptions: they cannot invert cryptographic functions unless explicitly allowed by the model. These symbolic constraints allow us to analyze potential attacks, such as replay attacks, man-in-the-middle attacks, or key compromise impersonation, within the Dolev-Yao framework.
In addition to the processes and messages, we specify the security properties of interest using queries, which ProVerif attempts to verify or refute. Typical properties include authentication, secrecy, and integrity. These properties are formally expressed in terms of reachability, correspondence assertions, or observational equivalence, depending on the specific security goals of the protocol under analysis.
By following this structured modeling methodology, we ensure that the symbolic model accurately reflects the intended behavior of the protocol, while accounting for the adversarial capabilities. This allows for the comprehensive verification of the security properties under consideration. Additionally, in order to ensure ProVerif is always able to complete a proof (or discover a counterexample), we ran our experiments on a research computing cluster with 88 cascade lake CPUs and 5TB of RAM.
@@ III-B Model Details
To construct our ProVerif models, we carefully and exhaustively analyzed the existing specifications of Olm [52], Megolm [51], and Sender Keys [50] [85], the canonical implementations of Olm and Megolm provided by the Matrix foundation [55] [54], the Matrix protocol specification [2], and much of the relevant online discussion [77]. We also closely compared our symbolic models with the existing computational constructions of Olm, Megolm, and Sender Keys [11] [17] to best ensure our models accurately capture the intended behaviors of the protocols. Additionally, although not central to our analysis, we construct Verifpal [63] models of of Olm, Megolm, and Sender Keys for the general benefit of the community.
Olm. We model the complete cryptographic handshake and the ratcheting mechanism of Olm. A diagram of the modeled handshake is shown in Appendix Section 1.
Megolm. We model the initialization of a Megolm session, the transmission of the ratchet key to peers in the Megolm session via Olm, the exchange of messages between peers in a Megolm session, and the incrementation of the Megolm ratchet. A diagram of the modeled cryptographic exchange is shown in Figure 2.
Signal. We employ a previously created Signal model, as presented in [62] and visualized in .
Sender Keys. We model the initialization of a Sender Keys session, the transmission of the ratchet key to peers in the Sender Keys session via Signal, the exchange of messages between peers in a Sender Keys session, and the incrementation of the Sender Keys ratchet. A diagram of the modeled cryptographic exchange is shown in Figure 2.
@@ III-C Abstractions and Limitations
Perfect Underlying Primitives. As in the symbolic cryptanalysis model, we assume the underlying cryptographic primitives are black-boxes. That is, we assume the primitives used in our models are secure and correct, particularly HKDF, HMAC, AEAD, and key signing. ProVerif and other cryptographic reasoning in tools in the symbolic model traditionally assume perfect computational cryptography.
Abstract Message Terms. Our models are abstract with respect to message formats, treating each message as a symbolic constructor. This abstract treatment of protocol messages is typical of symbolic tools.
No Ratchet Reseeding. The Megolm ratchet is designed to re-seed after every iterations as described in the Appendix VII-A. We do not capture this reseeding in our model — only forward iterations — as it is not relevant to the analysis under the Dolev-Yao model.
Finite Peers. We reason about a finite number of peers of a Megolm or Sender Keys group in our model, and prove our properties of interest under that assumption.
@@ IV-A Proven Properties of Interest
Using our ProVerif models, we automatically prove the Olm and Megolm properties as stated and motivated in section III-D. This includes Olm message confidentiality and authentication, Olm session key confidentiality, Olm forward secrecy (assuming Olm chooses to sign pre-keys), Olm post-compromise security, Megolm message confidentiality and authentication, and Megolm forward secrecy.
@@ IV-B3 Unknown Key-share in Megolm
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.3cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.1cm ]
[…]
@@ IV-C1 Olm identity key compromise without pre-key signing compromises the reliant Megolm session
no {msc}[ /msc/title top distance=0cm, /msc/first level height=.1cm, /msc/last level height=0.3cm, /msc/head height=0cm, /msc/instance width=0cm, /msc/head top distance=0.5cm, /msc/foot distance=-0.0cm, /msc/instance width=0cm, /msc/condition height=0.1cm ]
[…]
@@ V-A Computer-Aided Protocol Verification
In comparison, manual symbolic cryptanalysis has been theoretically feasible [29] but seldom employed until recent advances in proof-oriented programming languages. In particular, F*, a general-purpose proof-oriented language with support for dependent types proving, refinement times, executable semantics extractions, and verified implementation extractions to F#, C, or Wasm [81], has seen usage for this purpose. Additionally, other dependent types-based theorem provers such as Rocq [24] and Agda [69] are suited for this purpose, as they both offer executable semantics and implementation extractions 111Note, Rocq and Agda have different type theories: Rocq is based on the Calculus of Constructions (CoC) while Agda is based on Martin-Löf Type Theory. This difference results in different approaches to constructing proofs. It is entirely unexplored how this difference affects the formulation of cryptographic protocols via the methodology described in [29].. The theoretical framework for verifying protocols symbolically using dependent-types introduced in [29] has been formalized in F* and was used to analyze Signal [27] and later the Message Layer Security protocol [26].222Note that F* and dependent-types theorem provers are incredibly general and are expressive enough to capture all of mathematics. Interestingly, F* has been additionally used to construct verified cryptographic libraries [89] and to reason about group key agreement [25]. However, constructing such formalizations in a dependent types prover such as F* is generally much more time consuming than constructing models in a symbolic cryptanalysis tool.
@@ VI Conclusion
This paper has presented a comprehensive formal, symbolic analysis of the Olm and Megolm protocols within the Matrix cryptographic suite. Utilizing ProVerif, we constructed precise models and conducted automated symbolic analysis to verify key security properties, including confidentiality, authentication, forward secrecy, and post-compromise security for Olm and Megolm. Our approach enabled us to precisely and automatically reproduce known vulnerabilities and attacks.
In addition, through our comparative analysis with the Signal and Sender Keys protocols, we highlighted critical differences in post-compromise security, emphasizing the need for ongoing scrutiny and improvement of secure messaging protocols. We endorse the Matrix protocol specification’s current decision for Olm pre-keys to be signed, as our models confirm that when Olm pre-keys are signed the security guarantees of Olm and Megolm are comparable to Signal and Sender Keys.
As for future work, we plan to round out our symbolic analysis of the Matrix cryptographic suite with a formalization in the dependent types-based symbolic analysis framework DY*, which is implemented in F* [27]. This would allow us to construct secrecy and authentication proofs that are parameterized over the number of Olm ratchet iterations and Megolm epoch iterations, as well as extract executable semantics to more effectively reason about the implementations.
Overall, our findings underscore the importance of formal, mechanized proofs in ensuring the robustness of cryptographic systems, particularly in the context of federated and decentralized communication platforms like Matrix. We hope our work motivates further advancement of cryptographic verification tooling and the continued analysis and scrutiny of the protocols of which we all rely upon.